The area of a parellologram is given by the formula A=bh. The area of the parellogram is 66 sq units. The base is 2x+1. The height is x+1. Find its base and height.
Answer by Cromlix(4381)

Hi there,
A = bh
66 = (2x + 1)(x + 1)
66 = 2x^2 + 3x + 1
Collect like terms:
2x^2 + 3x - 65 = 0
(2x + 13)(x - 5)= 0
2x + 13 = 0
x = -13/2 (No answer as negative)
x - 5 = 0
x = 5
Base = 11 units
Height = 6 units
Hope this helps :-)